OpenStudy (anonymous):

f(xy)=f(x) and f(1) =3 then f(x) = ?

hartnn (hartnn):

put x = 1 in f(xy) = f(x) you will get f(y)

OpenStudy (anonymous):

will f(x) be 3? i.e a constant?

hartnn (hartnn):

yes

OpenStudy (anonymous):

hmm..fir kuch galat kia maine :/

OpenStudy (anonymous):

leme post whole q

hartnn (hartnn):

so this is not the original question?

OpenStudy (anonymous):

\[\LARGE \int\limits_1^{xy}f(t)dt=y \int\limits_1^{x}f(t)dt+x \int\limits_1^yf(t)dt\] for all x and y and y=R+,f(1)=3 then Options: f(x)=3log(3x) f(x)=log (base e)ex f'(y)=3/4 f'(x)=1/x

OpenStudy (anonymous):

\[\LARGE f(xy)(y+xy')=yf(x)+xf(y).y'\] y'=0 \[\LARGE yf(xy)=yf(x)\] \[\LARGE f(xy)=f(x)\] is it correct?

OpenStudy (anonymous):

@hartnn

OpenStudy (amistre64):

just tossing it around a little .... \[ \int_1^{xy}f(t)dt=y \int_1^{x}f(t)dt+x \int_1^yf(t)dt\] \[ F(xy)-F(1)=yF(x)-yF(1)+xF(y)-xF(1)\] \[ F(xy)-yF(x)-xF(y)=F(1)(1-y-x)\] \[ f(xy)(x'y+xy')-y'F(x)-yf(x)x'-x'F(y)-xf(y)y'=F(1)(1-y'-x')\]

OpenStudy (amistre64):

if this is wrt.x, then x'=1 \[f(xy)(y+xy')-y'F(x)-yf(x)-x'F(y)-xf(y)y'=-F(1)y'\] yeah, i cant see anyting from this this early :/

OpenStudy (zarkon):

Take the mixed partial with respect to x and y...simplify...you will get a simple equation in which you can solve for f

OpenStudy (anonymous):

what is wrong with my solution?